OpenStudy (anonymous):

Ms. Lyynch has 21 coins in nickel and dimes. Their total value is $1.65. How much of each coin does she have?

OpenStudy (anonymous):

x + y = 21 0.05x + 0.1y = 1.65 Can you solve this system of linear equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x + y = 21 x + 2y = 33 Now can you solve it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-x - y = -21 x + 2y = 33 ------------- y = 12 Now plug it back in and solve for x.

OpenStudy (anonymous):

oohh kay thxx

OpenStudy (anonymous):

what is y

OpenStudy (anonymous):

12.